OpenStudy (anonymous):

I'm terrible with time... Sasha finished mowing lawns at 4:15 PM. It took 1 3/4 h to mow the first lawn. The second and third lawns she mowed each took 1 1/5 h. She took one 45-min break. When did Sasha begin mowing the first lawn?

OpenStudy (unicornpoopcookies):

11:21

OpenStudy (unicornpoopcookies):

If anyone could help me with my question it would be very much appreciated,thanks.

OpenStudy (ajprincess):

The time she started mowing the thrd lawn can be found in this way. 1/5 of an hour=60/5mints=12mints 4.15 -1.12 ---- 3.03 ---- so at 3.03P.M she would have started mowing the lawn. following it?

OpenStudy (anonymous):

Took me a bit, but yeah, thanks again!

OpenStudy (ajprincess):

that's k:) Nw can u find the time at which started mowing the second lawn?

OpenStudy (ajprincess):

Did u solve for x? @UnicornPoopCookies

OpenStudy (anonymous):

3.03 -1.12 ------ 2.51?

OpenStudy (ajprincess):